OpenStudy (anonymous):

f(x)=√(x+8) and g(x)=4x a) domain of f(g(x)) b) domain of g(f(x))

OpenStudy (ghazi):

all the real numbers

OpenStudy (anonymous):

for both?

OpenStudy (ghazi):

yess i believe so

OpenStudy (ghazi):

let's see what @cwrw238 writes..she is more precised

OpenStudy (anonymous):

could you show the working out?

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

f(g(x)) = √(4x+8) now a square root of a negative number is not real so the domain is x >= -2 or [-2,+INF)

OpenStudy (anonymous):

is it -2 because -2 is the smallest negative number?

OpenStudy (ghazi):

\[f(g(x))= \sqrt{4x+8}=2\sqrt{x+2}\] and \[g(f(x))= 4\sqrt{x+8}\] so these are not defined for the values for what quantity under square root becomes negative, you've to exclude values x< -2, -8

OpenStudy (ghazi):

x< = -2,-8

OpenStudy (ghazi):

@Saeko did you get this?

OpenStudy (cwrw238):

its -2 because that gives sqrt(-8+8) = 0

OpenStudy (anonymous):

smaller than? wouldn't it be larger than? since it can't be a negative value

OpenStudy (anonymous):

oh i get it, except i'm sorta confused on the < or > part

OpenStudy (ghazi):

yes...for larger values it'll be positive

OpenStudy (anonymous):

so greater than?

OpenStudy (ghazi):

no...only less than...cuz for values greater than 2 function will hold the definition and it will be positive

OpenStudy (anonymous):

oooh...ok thanks :)

OpenStudy (ghazi):

:)